OpenStudy (anonymous):

Solve the following system of equations by the substitution method. 10x + 10y = 1 x = y - 3 What is the value of y? -29/20 31/20 -3/2

OpenStudy (anonymous):

Substitute x = y - 3 in the first equation and you should get : 10 ( y - 3 ) + 10y = 1 10y - 30 + 10 y = 1 Can you solve it now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, let's solve it further. 20y - 30 = 1 Now what do you do?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

Lol, still didn't get it? :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

20y = -29 So, \[y = \frac{ -29 }{ 20 }\]

OpenStudy (anonymous):

o ok thanks i get it now